OpenStudy (anonymous):

A rectangular pool is three times as long as it is wide. A redwood walkway 2 m wide with an area of 208 m^2 borders the pool. What are the dimensions of the pool?

OpenStudy (anonymous):

(3x + 2)(x + 2) = 3x^2 + 208 3x^2 + 8x + 4 = 3x^2 + 208 We can subtrate 3x^2 and 4 from both sides. 8x = 204 x = 25.5m width, 76.5m length (25.5)(76.5) = 1950.75 m^2 (27.5)(78.5) = 2158.75 m^2 The difference is 208m^2, this is a check.